Find an orthonormal basis of the plane in R3 defined by the equation 2a yz0
Answer #1

W = {(x, y, z) (2x+y+z=0} y=-2-ax (x, y, z) = (0, -2-2x, Z) = XC1, -2,0) +z (011) basis is {(1, -2,0), (0.-11)? then use Granالو. = V-U, V u.u - 60,-1,1)- (1, -20 = (-2/5, -15, 1) Ua 30 15 {es] is as onbonormal basis. basis. e- I TOfirst we find a basis for Plane W then we orthonormalize it using gramschmidth.